OpenStudy (anonymous):

the mass of 600 spherical lead pellets is found to be 66g their net weight when g=10ms-2 is a. 0.66n b.6.6n c.o.o66n d.66n

OpenStudy (anonymous):

mass , m = 66g* (1kg/1000g)=0.066kg || weight, w = m*g = 0.066*10 = 0.66N--------Answer ---choice a

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

yes thats correct one newton N is defined as the force which when applied to 1 kilogram mass gives it an acceleration of 1 metre per sec^2. f = m * a in this case the force is weight and acceleration is that due to gravity m = 0.066 kg and a = 10 ms^-2 = .066 * 10 = 0.66 N